OpenStudy (anonymous):

2e^x=30 round to 2nd decimal place

OpenStudy (asett):

What would you do first?

OpenStudy (princeharryyy):

e^x = 15 x*log(e) = log(15) x = 1.1760 x = 1.18

OpenStudy (anonymous):

im not sure

OpenStudy (princeharryyy):

@ineehel

OpenStudy (asett):

Nevermind : ). You would first divide each side by 2, to get e^x alone.

OpenStudy (asett):

Also, because e is a natural number, you would use the natural logarithm on your calculator: the 'ln' button, not the log button!

OpenStudy (anonymous):

so is the answer above correct?

OpenStudy (asett):

No. I will explain why.

OpenStudy (asett):

The 'log' button on your calculator means log base 10; however, the 'ln' button, or natural logarithm, basically means log base e (which is the natural number). Because of the two different bases, it does matter which one you use. In the question above, you can only use ln to cancel out the e, not the log. Does this make sense?

OpenStudy (anonymous):

2.71

OpenStudy (asett):

Whoo! Good job!

OpenStudy (anonymous):

thanks

OpenStudy (asett):

No problem!

OpenStudy (anonymous):

the easiest to solve this is is first divide by 2 to get e^x = 15, then the natural log of both side which will cancel the natural log on the left leaving you with x=ln(15). plug it into the calculator and voila...